Hello! I am working on my Algebra Homework and I am stuck on these 2 problems.. Can you please help me Tutor?
1) Give 3 equivalent fractions to 3/5. I could only think of 9/15!
2) Write in standard form: fifteen thousand, four and 53ten- thousandth.
Please Reply :) Thank You!
-Lauren

Answer by KnightOwlTutor(293)

1)3X2/5X2=6/10
3X4/5X4=12/20

2.15,453/10,1000
=1.5453